Given that log3 2= 0.631 what is the approximate value of log3 48
Answer by stanbon(75887)

Given that log3 2= 0.631 what is the approximate value of log3 48
---
log3(48) = log3(16)+log3(3)
---
= log3(2^4) + 1
---
= 4*log3(2) + 1
------
= 4*0.631 +1
------
= 3.524..
==========
Cheers,
Stan H.
===========